# Mass Translation and Reflectivity Device Solar Sail Momentum Management Using Predictive Control

### v2

@@ 3.2 Discretization - Mixed FOH and ZOH Tailored for Momentum Management Actuators
$\mbf{u}(\tau)={\bm{\Lambda}}^{-}_{k}(\tau)\mbf{u}_{k}+{\bm{\Lambda}}^{+}_{k}(\tau)\mbf{u}_{k+1}=\begin{bmatrix}\frac{t_{k+1}-\tau}{t_{k+1}-t_{k}}\mbf{1}&\mbf{0}\\ \mbf{0}&1\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{k}\\ u^{\text{RCD}}_{k}\end{bmatrix}+\begin{bmatrix}\frac{\tau-t_{k}}{t_{k+1}-t_{k}}\mbf{1}&\mbf{0}\\ \mbf{0}&0\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{k+1}\\ u^{\text{RCD}}_{k+1}\end{bmatrix}$. (18)
@@
where $\mbf{u}_{k}=\Big{[}\mbf{u}^{\text{AMT}^{\top}}_{k}\,\,\,u^{\text{RCD}}_{k}\Big{]}^{\top}$, $\mbf{A}_{k}=e^{\mbf{A}_{\ell}(t_{k+1}-t_{k})}$, $\mbf{B}_{w,k}=\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{w,\ell}\textrm{d}\tau$, $\mbf{B}_{u,k}^{-}=\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{u,\ell}{\bm{\Lambda}}^{-}_{k}(\tau)\textrm{d}\tau$, and $\mbf{B}_{u,k}^{+}=\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{u,\ell}{\bm{\Lambda}}^{+}_{k}(\tau)\textrm{d}\tau$. These LTV state-space matrices can be solved for as $\mbf{A}_{k}={\bm{\Phi}}_{x}(t_{k+1},t_{k})$, $\mbf{B}_{w,k}={\bm{\Phi}}_{w}(t_{k+1},t_{k})$, $\mbf{B}_{u,k}^{-}={\bm{\Phi}}_{u}^{-}(t_{k+1},t_{k})$, $\mbf{B}_{u,k}^{+}={\bm{\Phi}}_{u}^{+}(t_{k+1},t_{k})$, through the numerical integration of the matrix differential equations $\dot{{\bm{\Phi}}}_{x}(t,t_{k})=\mbf{A}_{\ell}{\bm{\Phi}}_{x}(t,t_{k}),$ $\dot{{\bm{\Phi}}}_{w}(t,t_{k})=\mbf{A}_{\ell}{\bm{\Phi}}_{w}(t,t_{k})+\mbf{B}_{w,\ell}$, $\dot{{\bm{\Phi}}}_{u}^{-}(t,t_{k})=\mbf{A}_{\ell}{\bm{\Phi}}_{u}^{-}(t,t_{k})+\mbf{B}_{u,\ell}{\bm{\Lambda}}^{-}_{k}$, and $\dot{{\bm{\Phi}}}_{u}^{+}(t,t_{k})=\mbf{A}_{\ell}{\bm{\Phi}}_{u}^{+}(t,t_{k})+\mbf{B}_{u,\ell}{\bm{\Lambda}}^{+}_{k}$, with initial values ${\bm{\Phi}}_{x}(t_{k},t_{k})=\mbf{1}_{n_{x}}$, ${\bm{\Phi}}_{w}(t_{k},t_{k})=\mbf{0}_{n_{x}\times n_{w}}$, ${\bm{\Phi}}_{u}^{-}(t_{k},t_{k})=\mbf{0}_{n_{x}\times n_{u}}$, and ${\bm{\Phi}}_{u}^{+}(t_{k},t_{k})=\mbf{0}_{n_{x}\times n_{u}}$ over the time interval $t\in[t_{k},t_{k+1}]$. The prediction model used by the proposed MPC algorithm is then chosen as the discrete-time LTV model in (19) evaluated at time $t_{k}$. Since the MPC prediction model accuracy significantly affects the performance, this approach is specifically tailored to account for the significantly different actuation natures of AMT and RCDs, where a FOH discretization is used to capture the smooth continuous translation of AMT and a ZOH discretization is used to model the discrete on-off pulsing of RCDs.
@@ 3.3 PWM-Based Quantization of Integer RCDs Torque and Actuation Thresholds
$t_{c}=\Delta t\cdot\frac{u^{\text{RCD}}_{k,\text{mpc}}}{u^{\text{RCD}}_{\text{max}}},$
@@
Given that frequent actuation, such as a large number of on-off RCD cycles, shortens actuator lifespan, reducing the number of non-zero control inputs is desirable. In this work, the amount of actuator usage is considered a performance metric for evaluating momentum management strategies. Inspired by the Solar Cruiser momentum management approach, where the actuator’s activation is governed by a threshold policy based on RW angular momentum bounds [11], a similar thresholding strategy is employed here. To avoid unnecessary short on-off RCD cycles, a threshold value $u_{\text{thr}}$ is chosen to define a deadband, where any continuous MPC input satisfying $|u^{\text{RCD}}_{k,\text{mpc}}|<u_{\text{thr}}$ is set to zero. The threshold value $u_{\text{thr}}$ acts as a tuning parameter to conserve control effort and minimize the on-off cycling of the RCDs, thereby reducing power consumption and prolonging actuator life. Fig. 4 illustrates an example of how the continuous MPC input is modified into a PWM pulse with the inclusion of a threshold at each momentum management time step.
@@ 3.4 Momentum Management Strategy 1: MPC with Single-Pulse PWM Quantized RCD Input
This section makes use of the results presented in Sections 3.1 through 3.3 to formulate the first proposed MPC policy for momentum management. The MPC policy is introduced first through the choices made with regards to the prediction model, state constraints, input constraints, and objective function used, followed a summary of the MPC policy in the form of an optimization problem.
@@ 3.4.2 State Constraints
$\mbf{h}_{b,\text{min}}^{\text{soft}}-{\bm{\alpha}}\leq\mbf{h}^{\text{RWs}}_{b,j|t_{k}}\leq\mbf{h}_{b,\text{max}}^{\text{soft}}+{\bm{\alpha}},$
where the soft constraint limits are given by $\mbf{h}_{b,\text{min}}^{\text{soft}}$ and $\mbf{h}_{b,\text{max}}^{\text{soft}}$, and $\bm{\alpha}$} is penalized in the MPC policy’s objective function in a quadratic fashion. This approach does not penalize the RWs’ angular momentum when $\mbf{h}_{b,\text{min}}^{\text{soft}}\leq\mbf{h}^{\text{RWs}}_{b,j|t_{k}}\leq\mbf{h}_{b,\text{max}}^{\text{soft}}$, as in this case the constraint is satisfied with ${\bm{\alpha}}=\mbf{0}$. Once these soft limits are exceeded, the soft constraint and quadratic penalty on $\bm{\alpha}$} helps keep the angular momentum away from the saturation limit.
@@ 3.4.3 Input Constraints
$\dot{\mbf{u}}^{\text{AMT}}_{\text{min}}\leq\frac{\mbf{u}^{\text{AMT}}_{j+1|t_{k}}-\mbf{u}^{\text{AMT}}_{j|t_{k}}}{t_{k+1}-t_{k}}\leq\dot{\mbf{u}}^{\text{AMT}}_{\text{max}}.$
@@ 3.4.4 Objective Function
$\sum_{j=0}^{N-1}\Bigl{(}\mbf{x}_{j|t_{k}}^{\top}\mbf{Q}\mbf{x}_{j|t_{k}}+\mbf{u}_{j|t_{k}}^{\top}\mbf{R}\mbf{u}_{j|t_{k}}+\tilde{\mbf{u}}^{\text{AMT}^{\top}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}\Bigl{)}+\mbf{x}_{N|t_{k}}^{\top}\mbf{P}\mbf{x}_{N|t_{k}}+\mbf{u}_{N|t_{k}}^{\top}\mbf{R}_{N}\mbf{u}_{N|t_{k}}+{\bm{\alpha}}^{\top}\mbf{C}{\bm{\alpha}}$
where the weights $\mbf{Q}=\mbf{Q}^{\top}$ and $\mbf{R}=\mbf{R}^{\top}$ are positive semi-definite and positive definite matrices, respectively, penalizing the state and control input over the prediction horizon of $N$ time steps. The terminal costs, $\mbf{P}$ and $\mbf{R}_{N}$, are associated with penalties on the state and input at the final step. The weight $\mbf{C}=\mbf{C}^{\top}$ is a positive semi-definite matrix used to adjust the emphasis on maintaining the soft constraint. The term $\tilde{\mbf{u}}^{\text{AMT}^{\top}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}$ within the objective function is defined with the variable $\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}=\mbf{u}^{\text{AMT}}_{j+1|t_{k}}-\mbf{u}^{\text{AMT}}_{j|t_{k}}$ and the positive definite weighting matrix ${\tilde{\mbf{R}}}={\tilde{\mbf{R}}}^{\top}$ is included to penalize translation of the AMT. This helps maintain the AMT at a stationary location, rather than allowing it to be in constant motion.
@@ 3.4.5 Summary of MPC Policy for Momentum Management Strategy 1
$\operatorname*{minimize}_{{\bm{\mathcal{X}}},\hskip 2.0pt{\bm{\mathcal{U}}},\hskip 2.0pt{\bm{\alpha}}}\sum_{j=0}^{N-1}\Bigl{(}\mbf{x}_{j|t_{k}}^{\top}\mbf{Q}\mbf{x}_{j|t_{k}}+\mbf{u}_{j|t_{k}}^{\top}\mbf{R}\mbf{u}_{j|t_{k}}+\tilde{\mbf{u}}^{\text{AMT}^{\top}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}\Bigl{)}+\mbf{x}_{N|t_{k}}^{\top}\mbf{P}\mbf{x}_{N|t_{k}}+\mbf{u}_{N|t_{k}}^{\top}\mbf{R}_{N}\mbf{u}_{N|t_{k}}+{\bm{\alpha}}^{\top}\mbf{C}{\bm{\alpha}}$ (21)
@@
The MPC policy in Section 3.4 provides a promising momentum management controller that can be implemented on a solar sail system. However, since the PWM-inspired quantization method discussed in Section 3.3 alters the optimal RCD input into a single pulse suitable for practical implementation, the MPC algorithm does not inherently account for this quantization when determining this optimal input, which affects its predictive capabilities. Incorporating knowledge of this quantization into the MPC framework has the potential to improve overall performance by aligning the prediction model closer to its practical implementation. The challenge with this is that the quantized RCD input is an integer variable, which transforms the MPC optimization problem into a mixed-integer optimization problem that cannot be solved efficiently and reliably. Typically, mixed-integer optimization problems require computationally-expensive branch-and-bound solution techniques [30, Chapter 8.3].
To address this, a framework to iteratively solve the mixed-integer MPC optimization problem as a sequence of QPs is proposed in this section. This iterative method quantizes the RCD input at the end of the prediction horizon step using a single PWM on-off pulse. The MPC optimization problem is then re-solved, treating this quantized input as a fixed virtual control input. With each iteration of the algorithm, the RCD input at the end of the shrinking horizon is fixed as a PWM-quantized input, while the remaining variables in the prediction horizon are optimized. Notably, only the RCD inputs are fixed during this process; state variables and AMT inputs remain as free design variables. The remainder of this section formulates the equations needed for this proposed iterative quantization strategy, followed by a summary of the proposed MPC policy.

### v3

@@ 4.2 Discretization - Mixed FOH and ZOH Tailored for Momentum Management Actuators
$\mbf{u}(\tau)$ $={\bm{\Lambda}}^{-}_{k}(\tau)\mbf{u}_{k}+{\bm{\Lambda}}^{+}_{k}(\tau)\mbf{u}_{k+1}$
$=\begin{bmatrix}\frac{t_{k+1}-\tau}{t_{k+1}-t_{k}}\mbf{1}&\mbf{0}\\ \mbf{0}&1\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{k}\\ u^{\text{RCD}}_{k}\end{bmatrix}+\begin{bmatrix}\frac{\tau-t_{k}}{t_{k+1}-t_{k}}\mbf{1}&\mbf{0}\\ \mbf{0}&0\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{k+1}\\ u^{\text{RCD}}_{k+1}\end{bmatrix}$, (28)
@@
where $\mbf{u}_{k}=\Big[\mbf{u}^{\text{AMT}^{\top}}_{k}\,\,\,u^{\text{RCD}}_{k}\Big]^{\top}$, $\mbf{A}_{k}=e^{\mbf{A}_{\ell}(t_{k+1}-t_{k})}$, $\mbf{B}_{w,k}=\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{w,\ell}\textrm{d}\tau$, $\mbf{B}_{u,k}^{-}=\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{u,\ell}{\bm{\Lambda}}^{-}_{k}(\tau)\textrm{d}\tau$, and $\mbf{B}_{u,k}^{+}=\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{u,\ell}{\bm{\Lambda}}^{+}_{k}(\tau)\textrm{d}\tau$. These LTV state-space matrices can be solved for as $\mbf{A}_{k}={\bm{\Phi}}_{x}(t_{k+1},t_{k})$, $\mbf{B}_{w,k}={\bm{\Phi}}_{w}(t_{k+1},t_{k})$, $\mbf{B}_{u,k}^{-}={\bm{\Phi}}_{u}^{-}(t_{k+1},t_{k})$, $\mbf{B}_{u,k}^{+}={\bm{\Phi}}_{u}^{+}(t_{k+1},t_{k})$, through the numerical integration of the matrix differential equations
$\dot{{\bm{\Phi}}}_{x}(t,t_{k})$ $=\mbf{A}_{\ell}{\bm{\Phi}}_{x}(t,t_{k}),$ (30)
$\dot{{\bm{\Phi}}}_{w}(t,t_{k})$ $=\mbf{A}_{\ell}{\bm{\Phi}}_{w}(t,t_{k})+\mbf{B}_{w,\ell},$ (31)
$\dot{{\bm{\Phi}}}_{u}^{-}(t,t_{k})$ $=\mbf{A}_{\ell}{\bm{\Phi}}_{u}^{-}(t,t_{k})+\mbf{B}_{u,\ell}{\bm{\Lambda}}^{-}_{k},$ (32)
$\dot{{\bm{\Phi}}}_{u}^{+}(t,t_{k})$ $=\mbf{A}_{\ell}{\bm{\Phi}}_{u}^{+}(t,t_{k})+\mbf{B}_{u,\ell}{\bm{\Lambda}}^{+}_{k}$, (33)
with initial values ${\bm{\Phi}}_{x}(t_{k},t_{k})=\mbf{1}_{n_{x}}$, ${\bm{\Phi}}_{w}(t_{k},t_{k})=\mbf{0}_{n_{x}\times n_{w}}$, ${\bm{\Phi}}_{u}^{-}(t_{k},t_{k})=\mbf{0}_{n_{x}\times n_{u}}$, and ${\bm{\Phi}}_{u}^{+}(t_{k},t_{k})=\mbf{0}_{n_{x}\times n_{u}}$ over the time interval $t\in[t_{k},t_{k+1}]$. The prediction model used by the proposed MPC algorithm is then chosen as the discrete-time LTV model in (29) evaluated at time $t_{k}$. Since the MPC prediction model accuracy significantly affects the performance, this approach is specifically tailored to account for the significantly different actuation natures of AMT and RCDs, where a FOH discretization is used to capture the smooth continuous translation of AMT and a ZOH discretization is used to model the discrete on-off pulsing of RCDs.
@@ 4.3 PWM-Based Quantization of Integer RCDs Torque and Actuation Thresholds
$t_{c}=\Delta t\cdot\frac{u^{\text{RCD}}_{k,\text{mpc}}}{u^{\text{RCD}}_{\text{max}}},$ (34)
@@
Given that frequent actuation, such as a large number of on-off RCD cycles, shortens actuator lifespan, reducing the number of non-zero control inputs is desirable. In this work, the amount of actuator usage is considered a performance metric for evaluating momentum management strategies. Inspired by the Solar Cruiser momentum management approach, where the actuator’s activation is governed by a threshold policy based on RW angular momentum bounds [20], a similar thresholding strategy is employed here. To avoid unnecessary short on-off RCD cycles, a threshold value $u_{\text{thr}}$ is chosen to define a deadband, where any continuous MPC input satisfying $|u^{\text{RCD}}_{k,\text{mpc}}|<u_{\text{thr}}$ is set to zero. The threshold value $u_{\text{thr}}$ acts as a tuning parameter to conserve control effort and minimize the on-off cycling of the RCDs, thereby reducing power consumption and prolonging actuator life. Figure 4 illustrates an example of how the continuous MPC input is modified into a PWM pulse with the inclusion of a threshold at each momentum management time step.
@@ 4.4 Momentum Management Strategy 1: MPC with Single-Pulse PWM Quantized RCD Input
This section makes use of the results presented in Sections 4.1 through 4.3 to formulate the first proposed MPC policy for momentum management. The MPC policy is introduced first through the choices made with regards to the prediction model, state constraints, input constraints, and objective function used, followed a summary of the MPC policy in the form of an optimization problem.
@@ 4.4.2 State Constraints
$\mbf{h}_{b,\text{min}}^{\text{soft}}-{\bm{\alpha}}\leq\mbf{h}^{\text{RWs}}_{b,j|t_{k}}\leq\mbf{h}_{b,\text{max}}^{\text{soft}}+{\bm{\alpha}},$ (36)
where the soft constraint limits are given by $\mbf{h}_{b,\text{min}}^{\text{soft}}$ and $\mbf{h}_{b,\text{max}}^{\text{soft}}$, and $\bm{\alpha}$} is penalized in the MPC policy’s objective function in a quadratic fashion. This approach does not penalize the RWs’ angular momentum when $\mbf{h}_{b,\text{min}}^{\text{soft}}\leq\mbf{h}^{\text{RWs}}_{b,j|t_{k}}\leq\mbf{h}_{b,\text{max}}^{\text{soft}}$, as in this case the constraint is satisfied with ${\bm{\alpha}}=\mbf{0}$. Once these soft limits are exceeded, the soft constraint and quadratic penalty on $\bm{\alpha}$} help keep the angular momentum away from the saturation limit. This allows MPC to preemptively react to momentum buildup before RW saturation, and relaxes the potential feasibility issue of using the saturation limit as a hard constraint. The values of $\mbf{h}^{\text{soft}}_{b,\text{min}}$ and $\mbf{h}^{\text{soft}}_{b,\text{max}}$ are selected based on the operation scenario and desired safety margin.
@@ 4.4.3 Input Constraints
$\dot{\mbf{u}}^{\text{AMT}}_{\text{min}}\leq\frac{\mbf{u}^{\text{AMT}}_{j+1|t_{k}}-\mbf{u}^{\text{AMT}}_{j|t_{k}}}{t_{k+1}-t_{k}}\leq\dot{\mbf{u}}^{\text{AMT}}_{\text{max}}.$ (37)
@@ 4.4.4 Objective Function
$\sum_{j=0}^{N-1}\Bigl(\mbf{x}_{j|t_{k}}^{\top}\mbf{Q}\mbf{x}_{j|t_{k}}+\mbf{u}_{j|t_{k}}^{\top}\mbf{R}\mbf{u}_{j|t_{k}}+\tilde{\mbf{u}}^{\text{AMT}^{\top}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}\Bigl)\\ +\mbf{x}_{N|t_{k}}^{\top}\mbf{P}\mbf{x}_{N|t_{k}}+\mbf{u}_{N|t_{k}}^{\top}\mbf{R}_{N}\mbf{u}_{N|t_{k}}+{\bm{\alpha}}^{\top}\mbf{C}{\bm{\alpha}}$ (38)
where the weights $\mbf{Q}=\mbf{Q}^{\top}$ and $\mbf{R}=\mbf{R}^{\top}$ are positive semi-definite and positive definite matrices, respectively, penalizing the state and control input over the prediction horizon of $N$ time steps. The terminal costs, $\mbf{P}$ and $\mbf{R}_{N}$, are associated with penalties on the state and input at the final step. The weight $\mbf{C}=\mbf{C}^{\top}$ is a positive semi-definite matrix used to adjust the emphasis on maintaining the soft constraint. The penalty ${\bm{\alpha}}^{\top}\mbf{C}{\bm{\alpha}}$ is activated only when soft constraint violation (${\bm{\alpha}}>\mbf{0}$). A higher value for $\mbf{C}$ places a stronger priority on minimizing the violation. The term $\tilde{\mbf{u}}^{\text{AMT}^{\top}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}$ within the objective function is defined with the variable $\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}=\mbf{u}^{\text{AMT}}_{j+1|t_{k}}-\mbf{u}^{\text{AMT}}_{j|t_{k}}$ and the positive definite weighting matrix ${\tilde{\mbf{R}}}={\tilde{\mbf{R}}}^{\top}$ is included to penalize translation of the AMT. This helps maintain the AMT at a stationary location, rather than allowing it to be in constant motion.
@@ 4.4.5 Summary of MPC Policy for Momentum Management Strategy 1
$\operatorname*{minimize}_{{\bm{\mathcal{X}}},\hskip 2.0pt{\bm{\mathcal{U}}},\hskip 2.0pt{\bm{\alpha}}}\sum_{j=0}^{N-1}\Bigl(\mbf{x}_{j|t_{k}}^{\top}\mbf{Q}\mbf{x}_{j|t_{k}}+\mbf{u}_{j|t_{k}}^{\top}\mbf{R}\mbf{u}_{j|t_{k}}+\tilde{\mbf{u}}^{\text{AMT}^{\top}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}\Bigl)$
$\hskip 150.00023pt+\mbf{x}_{N|t_{k}}^{\top}\mbf{P}\mbf{x}_{N|t_{k}}+\mbf{u}_{N|t_{k}}^{\top}\mbf{R}_{N}\mbf{u}_{N|t_{k}}+{\bm{\alpha}}^{\top}\mbf{C}{\bm{\alpha}}$ (39)
@@
$\hskip 20.00003pt\mbf{x}_{\text{min}}\leq\mbf{x}_{j|t_{k}}\leq\mbf{x}_{\text{max}},\hskip 11.49994ptj=0,\cdots,N,$
$\hskip 20.00003pt\mbf{u}_{\text{min}}\leq\mbf{u}_{j|t_{k}}\leq\mbf{u}_{\text{max}},\hskip 11.49994ptj=0,\cdots,N,$
$\hskip 20.00003pt\dot{\mbf{u}}^{\text{AMT}}_{\text{min}}\leq\frac{\mbf{u}^{\text{AMT}}_{j+1|t_{k}}-\mbf{u}^{\text{AMT}}_{j|t_{k}}}{t_{k+1}-t_{k}}\leq\dot{\mbf{u}}^{\text{AMT}}_{\text{max}},\hskip 10.00002ptj=0,\cdots,N-1,$
$\hskip 20.00003pt\mbf{h}_{b,\text{min}}^{\text{soft}}-{\bm{\alpha}}\leq\mbf{h}^{\text{RWs}}_{b,j|t_{k}}\leq\mbf{h}_{b,\text{max}}^{\text{soft}}+{\bm{\alpha}},\hskip 11.49994ptj=0,\cdots,N,$
@@
The MPC policy in Section 4.4 provides a promising momentum management controller that can be implemented on a solar sail system. However, since the PWM-inspired quantization method discussed in Section 4.3 alters the optimal RCD input into a single pulse suitable for practical implementation, the MPC algorithm does not inherently account for this quantization when determining this optimal input, which affects its predictive capabilities. Incorporating knowledge of this quantization into the MPC framework has the potential to improve overall performance by aligning the prediction model closer to its practical implementation. The challenge with this is that the quantized RCD input is an integer variable, which transforms the MPC optimization problem into a mixed-integer optimization problem that cannot be solved efficiently and reliably. Typically, mixed-integer optimization problems require computationally-expensive branch-and-bound solution techniques [40, Chapter 8.3].
To address this, a framework to iteratively solve the mixed-integer MPC optimization problem as a sequence of QPs is proposed in this section. This iterative method quantizes the RCD input at the end of the prediction horizon step using a single PWM on-off pulse. The MPC optimization problem is then re-solved, treating this quantized input as a fixed virtual control input. With each iteration of the algorithm, the RCD input at the end of the shrinking horizon is fixed as a PWM-quantized input, while the remaining variables in the prediction horizon are optimized. Notably, only the RCD inputs are fixed during this process; state variables and AMT inputs remain as free design variables. The remainder of this section formulates the equations needed for this proposed iterative quantization strategy, followed by a summary of the MPC policy.